Grade 11General Physics

The specific gravity of seawater is 1.03 and that of ice is 0.92 .What fraction of an ice berg is above the surface of water?

Sayed Munazza Saniya
8 Years agoGrade 11

1 Answer

Arun
8 Years ago
Dear student
 
Let mass of iceberg be m. In terms of volume V, m = V*0.92 

Let portion immersed be of volume V1, then it also displaces the same VOLUME of seawater. By Archimedes Principle, the buoyance force is density of seawater * V1. 

At equilibrium 
mg = 1.03*V1*g 
0.92*V*g = 1.03*V1*g 
So, V1/V = 0.92/1.03 = 0.8932

Volume above seawater = 0.106 of total VOLUME
 
 
 
Regards
Arun (askIITians forum expert)